Determine the max. bending stress in the beam. Shear stress at point H, max. shear stress

Psky loom 5 b) 3m - 75 у .
Answer #1

SOLUTION : 8 KN Tim. RA 3 197. RB www . me {MA = 0 ; EFvEO. 8 E RA TRB. 40. EMAZO. 8 (1.9) = RB (3), 1 RB = 5.0667 KN. 0 Đ |The Mmasi = 2:9333xloq Mman. TMmon & 5.573 kNm) BMD ) Mmare = 5573.33 Nm Scanned with CamScannerOp = max sman. 02 25X751 1- 100 75 X 100 men 12 / 1 1 - 6250000 - 1757812:5 I = 4492187.55 mm - A4, + A292 mon = ATA2 ImaxOb man = 77541982.61 N/m2 Jobman - 97.541 Mpal (ii) 7 Case (i at point H. suppose It lies just - web above. portion.. --------- - 37:5- 125 = 25 mm at @ Flange flange -BX103x_93715 x 25 = 44921875x75 0.556 Mpa. H.@ Flange.--- H @ Tease (ii) H Ties juBX10° x 93115X25 H@ web portion ; 44.92187.55 x 25 | ZH @ web = 1668 MPa (iii) e marr -62.5 B A = 62. 5x25 A = 15.62.5 mm? ýman = 3.478 Meal --- - -------- ------- ---- - RESULTS (1) bona = 77.541 MPa. (11) The Flange = 0.556 MPa- a web = 1.668 MPa.